please solve.
120- 3/4 y = 60

Answer by stanbon(75887)

120- 3/4 y = 60
(3/4)y=120-60
(3/4)y=60
Multiply both sides by (4/3) to get:
y=60(4/3)
y=80
Cheers,
Stan H.